# Vision-Language Models for Medical Report Generation and Visual Question Answering: A Review

### v1

@@ 2.2 Natural Language Processing
In recent years, there has been a remarkable advancement in NLP mainly due to the development of the Transformer [Vas+17]. Beyond incorporating embedding layers and positional encodings, the Transformer architecture consists of an encoder that processes input data, represented by vectors obtained from embedded and positionally encoded tokens. The encoder-generated representation then serves as the input for the subsequent decoder, which transforms these vector representations into a relevant output tailored to the specific task at hand. A defining characteristic of the Transformer lies in its self-attention mechanism, notably the scaled dot-product attention, which proves instrumental in capturing intricate dependencies within sequences. This mechanism, employed in both the encoder and decoder, utilizes queries and keys during the attention process. Queries serve as projections of the input sequence, encapsulating information for attending to other positions, while keys represent the positions within the sequence. Enhanced by multi-head attention for parallelization, the self-attention mechanism enables the model to dynamically weigh different parts of the input sequence, fostering a nuanced understanding of contextual relationships. The incorporation of potent GPUs significantly amplifies these capabilities, ensuring swift and efficient computations. Furthermore, each layer in both the encoder and decoder encompasses sub-layers, including a feedforward NN, further augmenting the model’s capacity to capture intricate patterns within the data. In practice, Transformers face limitations in effectively processing long sequences and exhibit less selectivity about relevant information when considering all positions in the sequence. To address these issues, various techniques have been proposed. One such approach, known as hierarchical attention [Yan+16], strategically reduces computational complexity and enhances contextual sensitivity, by initially computing attention at the word level, and then at the sentence level. Another notable advancement in attention algorithms is the FlashAttention-2 [Dao23], designed to significantly accelerate attention computations.
@@
The synergy between the enhanced computational power provided by GPUs and the advancements in attention mechanisms has played a pivotal role in the development of large language models (LLMs). These models are meticulously trained on vast datasets, accumulating a substantial number of parameters, including weights and biases within the neural network. Initial LLMs include, but are not limited to, BERT (Bidirectional Encoder Representations from Transformers) [Dev+19] (the largest version comprising $235$ M parameters), ALBERT (A Lite BERT) [Lan+19] (with the largest variant at $12$ M parameters), and Megatron-LM [Sho+19] (the largest version featuring $1.2$ B parameters). The era of even larger LLMs began in $2020$, introducing models like GPT-3 (the 3-generation Generative Pre-trained Transformer) [Bro+20] ($175$ B parameters) and PaLM (Pathways Language Model) [Cho+22] ($540$ B parameters). Some of the most recent LLMs are LLaMA (Large Language Model Meta AI) [Tou+23a], Vicuna [Chi+23], Llama 2 [Tou+23], and Mistral [Jia+23]. Note that LLMs can also be used to generate word embeddings (e.g., BERT [Dev+19]).
2.3 Computer Vision
@@
Computer vision (CV) involves interpreting and understanding the world from their images or videos [Ji20]. Data in CV is encoded as numerical values representing the intensity or brightness of pixels. The extraction of visual patterns like edges, textures, and objects in images or video frames serves as building blocks for various CV tasks. Image classification is the task of assigning a label to an entire image, determining the main object or scene. Object detection involves identifying and locating multiple objects within an image, providing both labels and bounding boxes. Image segmentation divides an image into meaningful segments, assigning a label to each pixel and outlining the boundaries of distinct objects or regions. Various machine learning techniques and models are utilized for these tasks [Mah+22].
@@
The Convolutional Neural Networks (CNNs) [Yam+18] represents a significant advancement in CV, especially in image classification. Besides pooling and fully connected layers, CNNs have also convolution layers, which are instrumental in applying convolution operations. During a convolution operation, a small filter or kernel slides over the input data. At each position, the filter performs element-wise multiplications with the local regions of the input. The results of these multiplications are then summed, creating a new value in the output feature map. This process is repeated across the entire input, capturing patterns and features at different spatial locations. A commonly used CNN architecture is ResNet (Residual Network), or RN [He+16], which diverges from traditional NNs by focusing on learning the differences (residuals) between desired and actual outputs at each layer. This innovative approach aids in training exceptionally deep NNs by mitigating the vanishing gradient problem, providing shortcuts for more effective gradient flow throughout the network. Another popular CNN is DenseNet (Dense Convolutional Network) [Hua+19]. In its architecture, each layer receives feature maps from all preceding layers as input, and it, in turn, provides its feature maps as input to all subsequent layers. This dense connectivity pattern promotes feature reuse, and parameter efficiency, and enhances the network’s ability to capture intricate patterns across different scales. EfficientNet [TL20] represents a CNN architecture with a unique scaling method that uniformly scales network width, depth, and resolution in a balanced manner. This approach allows the network to achieve improved performance while maintaining computational efficiency
A pivotal advancement in object detection has been the emergence of R-CNNs (region-based CNNs) [Gir+14]. The primary objective of R-CNNs (e.g., faster R-CNN [Ren+15]) is to not only categorize objects but also precisely determine their spatial location in an image with a bounding box. Contrary to traditional CNNs, which analyze the entire image in a single pass, R-CNNs utilize a three-step approach. First, potential regions of interest, called region proposals, are extracted (e.g., using a selective search or a region proposal network). Next, each region is fed into a CNN, and then the output feature vectors are classified (e.g., by the support vector regression (SVM)). R-CNN models show a high accuracy performance, especially for larger objects. Another known object detection model is YOLO (You Only Look Once) [Red+16], which divides an image into a grid and predicts bounding boxes and class probabilities for each grid cell in a single forward pass. This approach enables the model to detect multiple objects in an image simultaneously with higher speed which makes it a popular choice for real-time applications.
The CNNs have also revolutionized image segmentation. Mask R-CNN [He+17], an extension of faster R-CNN [Ren+15], introduces an additional branch dedicated to outlining precise pixel-level segmentation for each object instance. U-Net [RFB15], featuring a U-shaped structure, proves especially popular in biomedical fields. The architecture is characterized by a contracting path responsible for capturing context and extracting features from the input image (encoding) and an expansive path that facilitates precise localization and segmentation of objects (decoding). Moreover, Fully Convolutional Networks (FCNs) (e.g., FastFCN [Wu+19]), represent a category of CNNs where traditional fully connected layers are replaced with convolutional layers, making them exceptionally well-suited for image segmentation tasks.
@@
Transformer models which were originally designed for NLP tasks, have also found valuable applications in CV. For instance, vision transformers (ViTs) [Dos+21] can capture intricate relationships and dependencies across the entire image. This is achieved by leveraging the Transformer architecture and treating images as sequences of smaller patches. Here, each image patch undergoes a process of flattening into a vector, followed by passage through an embedding layer. The embedding layer enriches the flattened image patches, providing a more expressive and continuous representation. Next, positional encodings are incorporated into the embeddings, conveying information about the spatial arrangement of the image patches. A distinctive feature of ViTs is the introduction of a special token designed to capture global information about the entire image. This special token has an associated learnable token embedding, represented by a vector with its unique set of parameters. ViTs have achieved notable success in segmentation [RBK21], anomaly detection [Mis+21], medical image classification [Man+23] and even outperformed CNNs in some cases [Tya+21, Xin+22].
3 Background on Vision-Language Models
@@
Many real-world situations inherently involve a variety of data modalities. For example, autonomous cars must process information from various sensors like cameras, RADAR, and GPS to ensure safe and effective navigation [Par+22]. Similarly, in oncology, the amalgamation of radiology images with genomic data, histopathology images, and clinical reports has the potential to greatly refine the precision of cancer diagnoses [Boe+21]. This motivated the development of VLMs, which can handle and understand both NLP and CV data simultaneously.
@@ 3.1 Model Architecture
Based on how different modalities are fused in VLMs, they are generally categorized into two groups [Che+23]: (1) single-stream (e.g., VisualBERT [Li+19], UNITER (UNiversal Image-TExt Representation Learning) [Che+19]), and (2) dual-stream models (e.g., ViLBERT (Vision-and-Language BERT) [Lu+19], CLIP [Rad+21]).
@@
A single-stream VLM adopts an efficient architecture for processing both visual and textual information within a unified module. This architecture incorporates an early fusion of distinct modalities, where feature vectors from various sources are concatenated into a single vector (e.g., MedViLL [Moo+22]). Subsequently, this combined representation is fed into a single stream. One notable advantage of the single-stream design is its parameter efficiency, achieved by employing the same set of parameters for all modalities. This not only simplifies the model but also contributes to computational efficiency during both training and inference phases [Che+23].
@@
In contrast, a dual-stream VLM extracts visual and textual representations separately in parallel streams that do not share parameters. This architecture usually has higher computationally complex compared to single-stream architectures. Visual feature vectors are generated from pre-trained vision encoders, such as CNNs, R-CNNs, or ViTs, and textual feature vectors are obtained from pre-trained text encoders, usually based on a Transformer architecture (e.g., PubMedCLIP [EMD23]). They are then inputted into a multimodal fusion module, often leveraging attention mechanisms, to integrate information from both modalities and to create cross-modal representations. This late fusion approach allows for more intricate interactions between visual and textual information, enabling the model to capture complex cross-modal dependencies. However, it comes at the cost of increased computational complexity compared to single-stream architecture.
@@
In addition, modality-combined representations may be optionally forwarded into a decoder before producing the final output. Consequently, VLMs are classified into two groups: (1) encoder-only (e.g., ALIGN (A Large-scale ImaGe and Noisy-text embedding) [Jia+21]), and (2) encoder-decoder models (e.g., SimVLM (Simple Visual Language Model) [Wan+22b]). Encoder-only models are advantageous in scenarios where the primary objective is efficient representation learning. They often exhibit streamlined processing and reduced computational complexity, making them suitable for tasks requiring compact and informative representations. However, these models might lack the capability to generate intricate and detailed outputs, limiting their use in tasks demanding nuanced responses or creative generation. On the other hand, encoder-decoder models offer the flexibility to generate complex and diverse outputs, making them well-suited for tasks like image captioning, translation, or any application requiring creative responses. The decoding step allows for the transformation of joint representations into meaningful outputs. However, this versatility comes at the cost of increased computational load and complexity.
@@ 3.2 Model Training
In this section, we discuss strategies for adapting VLMs through full or partial model updates.
@@ 3.2.1 Transfer Learning
In the realm of VLMs and the broader scope of machine learning, a commonly adopted strategy is to employ pre-trained models and customize them to specific downstream tasks, a method commonly known as transfer learning. This process typically involves fine-tuning the model’s parameters using smaller task-specific datasets to address the intricacies of the target task. Moreover, during fine-tuning, one may adjust the model’s architecture. This can include modifications to the final layers or the introduction of new layers, such as classification or regression layers, tailored to meet the specific requirements of the task at hand. The underlying idea is to adapt the pre-trained model to the specifics of the new task while retaining the knowledge it gained during the initial pre-training.
@@ 3.2.2 Curriculum Learning
Curriculum learning [Sov+21] presents an innovative approach when dealing with tasks or data exhibiting a natural progression or hierarchy. This method involves strategically presenting training examples or tasks in a designed order, typically based on difficulty or complexity measures. The recent VLM, LLaVa-Med [Li+23], adopts curriculum learning during its training phase. This allows the model to learn gradually, starting with simpler examples and progressing to more intricate ones. This orchestrated learning sequence enhances the model’s adaptability and performance.
3.2.3 Pre-Training Tasks
@@
One prevalent task is contrastive learning (CL), a technique that encourages the model to learn meaningful representations by contrasting positive pairs with negative pairs of both visual and textual data [Li+21]. In contrastive learning, the model is trained to map both positive and negative pairs into a shared embedding space. Positive pairs consist of examples where the visual and textual content are related, such as an image paired with its corresponding textual description. Conversely, negative pairs consist of examples where the visual and textual content are unrelated, like an image paired with a randomly selected different textual description. The objective is to maximize the distance between positive pairs while maximizing the distance between negative pairs in the shared embedding space. Several contrastive loss functions are used in contrastive learning to achieve this objective. Various contrastive loss functions are employed to achieve this objective, with the InfoNCE (Noise-Contrastive Estimation) loss [OLV19] being a common choice. InfoNCE formulates a probabilistic task where the model is trained to maximize the likelihood of observing positive pairs and minimize the likelihood of observing negative pairs. The negative log-likelihood of the positive pair is used as the loss. The CLIP [Rad+21] employs InfoNCE loss with cosine similarity. Other VLMs deviate from InfoNCE, for instance, ALIGN [Jia+21] uses a normalized softmax loss. This loss computes the softmax over cosine similarities between the normalized embeddings of positive and negative pairs, aiming to boost positive similarity while diminishing negative similarities.
@@
In NLP pre-training, a widely used task is masked language modeling (MLM) [Tay53]. It was first introduced and applied in the BERT model [Dev+19]. MLM involves randomly selecting a percentage of tokens within textual data and replacing them with a special token, often denoted as MASK. The model predicts these masked tokens by taking into account the context on both sides of them, allowing the model to grasp nuanced contextual information. VLMs such as UNITER [Che+19] and VisualBERT [Li+19] leverage MLM for pre-training. Extending this idea to images gave rise to masked image modeling (MIM) [Xie+22]. Within MIM, certain patches are masked, prompting the model to predict the content of masked regions. This process enables the model to draw context from the entirety of the image, encouraging the integration of both local and global visual features. VLMs like UNITER [Che+19] and ViLBERT [Lu+19] leverage MIM for enhanced performance. The cross-entropy loss is employed in MLM and MIM tasks to measure the difference between predicted and actual probability distributions for the masked elements. Additionally, MLM can be combined with MIM, allowing the reconstruction of the masked signal in one modality with support from another modality [Kwo+23].
Self-supervised learning [Ran+23] is a fundamental paradigm in training VLMs, offering a powerful alternative to traditional supervised learning by allowing models to generate their own labels from the data. This is particularly beneficial when obtaining large amounts of labeled data is challenging or expensive. In self-supervised learning for VLMs, the models formulate tasks that leverage inherent structures within the data, enabling them to learn meaningful representations across modalities without explicit external labels. CL, MLM, and MIM are examples of self-supervised learning tasks.
[…]
@@
In VLM pre-training, multiple tasks are often combined in a unified framework, allowing models to grasp nuanced contextual information across modalities. The final loss function can combine contrastive loss, cross-entropy loss for masked token prediction, and other task-specific losses. This comprehensive pre-training approach equips VLMs with versatile representations for diverse downstream tasks. For example, ALBEF [Li+21] adopts a comprehensive pre-training objective encompassing three tasks: CL, MLM, and ITM. The overall loss is then computed as the sum of these individual components.

### v2

@@ 2.2.4 Transformers
In recent years, there has been a remarkable advancement in NLP mainly due to the development of the Transformer models [vaswani2017attention]. Beyond incorporating embeddings and positional encodings, the Transformer architecture consists of an encoder that processes input data, represented by vectors obtained from embedded and positionally encoded tokens. The encoder-generated representation then serves as the input for the subsequent decoder, which transforms these vector representations into a relevant output tailored to the specific task at hand. A defining characteristic of the Transformer lies in its self-attention mechanism, notably the scaled dot-product attention, which proves instrumental in capturing intricate dependencies within sequences. This mechanism, employed in both the encoder and decoder, utilizes queries and keys during the attention process. Queries serve as projections of the input sequence, encapsulating information for attending to other positions, while keys represent the positions within the sequence. Enhanced by multi-head attention for parallelization, the self-attention mechanism enables the model to dynamically weigh different parts of the input sequence, fostering a nuanced understanding of contextual relationships. Each layer in both the encoder and decoder encompasses sub-layers, including a feedforward NN, further augmenting the model’s capacity to capture intricate patterns within the data. In practice, Transformers face limitations in effectively processing long sequences and exhibit less selectivity about relevant information when considering all positions in the sequence. Various techniques have been proposed to address these issues. One such approach, known as hierarchical attention [yang2016hierarchical], strategically reduces computational complexity and enhances contextual sensitivity by initially computing attention at the word level and then at the sentence level. Another notable advancement in attention algorithms is the FlashAttention [dao2022flashattention] and FlashAttention-2 [dao2023flashattention2], designed to accelerate attention computations significantly.
@@
The synergy between the enhanced computational power provided by Graphical Processing Units (GPUs) and the advancements in attention mechanisms has played a pivotal role in the development of large language models (LLMs). These models are meticulously trained on vast datasets with a very large number of parameters. Initial LLMs include but are not limited to, BERT (Bidirectional Encoder Representations from Transformers) [devlin2019bert] (the largest version comprising $235$ M parameters), ALBERT (A Lite BERT) [lan2019albert] (with the largest variant at $12$ M parameters), and Megatron-LM [shoeybi2020megatronlm] (the largest version featuring $1.2$ B parameters). The era of even larger LLMs began in $2020$, introducing models like GPT-3 (the 3-generation Generative Pre-trained Transformer) [brown2020gbt] ($175$ B parameters) and PaLM (Pathways Language Model) [chowdhery2022palm] ($540$ B parameters). Some of the most recent LLMs are LLaMA (Large Language Model Meta AI) [touvron2023llama], Vicuna [chiang2023vicuna], Llama 2 [touvron2023llama2], and Mistral [jiang2023mistral]. Note that encoder-only LLMs can be used to generate token embeddings (e.g., BERT [devlin2019bert] or GatorTron [yang2022gatortron]).
2.3 Computer Vision (CV)
@@
CV involves interpreting and understanding the world from their images or videos [ji2020cv]. Data in CV is encoded as numerical values representing the intensity or brightness of pixels. The extraction of visual patterns like edges, textures, and objects in images or video frames serves as building blocks for various CV tasks. Image classification is the task of assigning a label to an entire image, determining the main object or scene. Object detection involves identifying and locating multiple objects within an image, providing both labels and bounding boxes. Image segmentation divides an image into meaningful segments, assigning a label to each pixel and outlining the boundaries of distinct objects or regions. Various ML techniques and models are utilized for these tasks [mahadevka2022MLinCV].
2.3.1 Convolutional Neural Networks (CNNs)
@@
CNNs represent a significant advancement in CV [yamashita2018cnn]. Besides pooling and fully connected layers, CNNs also have convolution layers, which apply convolution operations to input data. During a convolution operation, a small filter or kernel slides over the input data. At each position, the filter performs element-wise multiplications with the local regions of the input. The results of these multiplications are then summed, creating a new value in the output feature map. This process is repeated across the entire input, capturing patterns and features at different spatial locations. The well-known CNNs include Residual Network (ResNet) [he2016resnet], Dense Convolutional Network (DenseNet) [huang2019densenet], Efficient Network (EfficientNet) [tan2020efficientnet] and many others.
2.3.2 Vision Transformers (ViTs)
@@
Transformer models, which were originally proposed for NLP tasks, have also found valuable applications in CV. For instance, the ViT model [dosovitskiy2021vit] can capture intricate relationships and dependencies across the entire image. This is achieved by leveraging the Transformer architecture and treating images as sequences of smaller patches. Each image patch undergoes a process of flattening into a vector, followed by passage through an embedding layer. The embedding layer enriches the flattened image patches, providing a more expressive and continuous representation. Next, positional encodings are incorporated into the embeddings, conveying information about the spatial arrangement of the image patches. A distinctive feature of ViTs is the introduction of a special token designed to capture global information about the entire image. This special token has an associated learnable token embedding, represented by a vector with its unique set of parameters. ViTs have achieved notable success in semantic segmentation [ranftl2021visiontf], anomaly detection [mishra2021vtadl], medical image classification [manzari2023medvit] and even outperformed CNNs in some cases [tyagil2021detpneum, xin2022transfskin].
3 Vision-Language Models (VLMs)
@@
Many real-world situations inherently involve a variety of data modalities. For example, autonomous cars must process information from various sensors like cameras, RADAR, LiDAR, and/or GPS to ensure safe and effective navigation [parekh2022autocars]. Similarly, in cancer care, the fusion of radiology images with genomic data, digitized histopathology slides, and clinical reports has the potential to improve diagnosis, treatment planning, and post-treatment surveillance [boehm2021advanceoncology, waqas2023multimodaloncology, Mohsan2023]. This motivated the development of VLMs, which can handle and understand both NLP and CV data simultaneously.
@@ 3.1 Model Architecture
Based on how different data modalities are fused together in VLMs, they are generally categorized into two groups [chen2022vlp]: (1) single-stream (e.g., VisualBERT [li2019visualbert] and UNITER or UNiversal Image-TExt Representation Learning [chen2019uniter]), and (2) dual-stream models (e.g., ViLBERT Vision-and-Language BERT [lu2019vilbert] and CLIP or Contrastive Language-Image Pre-training [radford2021clip]).
Single-Stream Models
@@
A single-stream VLM adopts an efficient architecture for processing both visual and textual information within a unified module. This architecture incorporates an early fusion of distinct data modalities, where feature vectors from various data sources are concatenated into a single vector (e.g., MedViLL [moon2022medvill]). Subsequently, this combined representation is fed into a single stream. One notable advantage of the single-stream design is its parameter efficiency, achieved by employing the same set of parameters for all modalities. This not only simplifies the model but also contributes to computational efficiency during both training and inference phases [chen2022vlp].
Dual-Stream Models
@@
A dual-stream VLM extracts visual and textual representations separately in parallel streams that do not share parameters. This architecture usually has higher computational complexity than single-stream architectures. Visual features are generated from pre-trained vision encoders, such as CNNs or ViTs, and textual features are obtained from pre-trained text encoders, usually based on a Transformer architecture (e.g., PubMedCLIP [eslami2023pubmedclip]). Both features are then fed into a multimodal fusion module, often leveraging attention mechanisms, to integrate information from both data modalities and to learn cross-modal representations. This late fusion approach allows for more intricate interactions between visual and textual information, enabling the model to capture complex cross-modal dependencies. However, it comes at the cost of increased computational complexity compared to single-stream architecture.
3.1.2 Encoder vs. Encoder-Decoder VLMs
The learned cross-modal representations can be optionally processed by a decoder before producing the final output. Consequently, VLMs are classified into two groups: (1) encoder-only (e.g., ALIGN (A Large-scale ImaGe and Noisy-text embedding) [jia2021align]), and (2) encoder-decoder models (e.g., SimVLM (Simple Visual Language Model) [wang2022simvlm]).
Encoder-only Models
@@
These models are advantageous in scenarios where the primary objective is efficient representation learning. They often exhibit streamlined processing and reduced computational complexity, making them suitable for tasks requiring compact and informative representations. However, these models might lack the capability to generate intricate and detailed outputs, limiting their use in tasks demanding nuanced responses or creative generation.
Encoder-Decoder Models
These models offer the flexibility to generate complex and diverse outputs, making them well-suited for tasks like image captioning, translation, or any application requiring creative responses. The decoding step allows for the transformation of joint representations into meaningful outputs. However, this versatility comes at the cost of increased computational load and complexity.
@@ 3.2.1 Transfer Learning
A commonly adopted strategy in ML is to employ pre-trained models and customize them to specific downstream tasks — a method commonly known as transfer learning. This process typically involves fine-tuning the model’s parameters using smaller task-specific datasets to address the intricacies of the target task [bommasani2022opportunities]. Transfer learning can also be considered as the process of starting the parameters optimization for a task using a set of already optimized parameters on another task instead of using random initialization. Transfer learning may involve some modification of the original model’s architecture. This can include modifications to the final layers or the introduction of new layers, such as classification or regression layers, tailored to meet the specific requirements of the task at hand [bommasani2022opportunities]. The underlying idea is to adapt the pre-trained model to the specifics of the new task while retaining the knowledge it gained during the initial pre-training.
@@ 3.2.2 Curriculum Learning
Curriculum learning presents an innovative approach when dealing with tasks or data exhibiting a natural progression or hierarchy. This method involves strategically presenting training examples or tasks in a designed order, typically based on difficulty or complexity measures [soviany2021currlearn]. The recent medical VLM, LLaVa-Med [li2023llavamed], adopts curriculum learning during its training phase. This allows the model to learn gradually, starting with simpler examples and progressing to more intricate ones. This orchestrated learning sequence enhances the model’s adaptability and performance.
3.2.3 Self-Supervised Learning (SSL)
SSL is a fundamental paradigm in training VLMs, offering a powerful alternative to traditional supervised learning by allowing models to generate their own labels from the data [rani2023selfsupervised]. This is particularly beneficial when obtaining large amounts of labeled data is challenging or expensive. In self-supervised learning for VLMs, the models formulate tasks that leverage inherent structures within the data, enabling them to learn meaningful representations across modalities without explicit external labels. Contrastive learning, masked language modeling, and masked image modeling (described in the following sub-section) are examples of self-supervised learning tasks.
3.2.4 Pre-Training Process and Tasks
@@
CL encourages the model to learn meaningful representations by contrasting positive pairs with negative pairs of both visual and textual data [li2021albef]. During CL, the model is trained to map both positive and negative pairs into a shared embedding space. Positive pairs consist of examples where the visual and textual content are related, such as an image paired with its corresponding textual description. Conversely, negative pairs consist of examples where the visual and textual content are unrelated, like an image paired with a randomly selected different textual description. The objective is to bring positive pairs closer together while pushing negative pairs farther apart in the shared embedding space. Various contrastive loss functions are employed to achieve this objective, with the InfoNCE (Noise-Contrastive Estimation) loss [oord2019infonceloss] being a common choice. InfoNCE formulates a probabilistic task where the model is trained to maximize the likelihood of observing positive pairs and minimize the likelihood of observing negative pairs. The negative log-likelihood of the positive pair is used as the loss. CLIP [radford2021clip] employs InfoNCE loss with cosine similarity. On the other hand, ALIGN [jia2021align] uses a normalized softmax loss. This loss computes the softmax over cosine similarities between the normalized embeddings of positive and negative pairs, aiming to boost positive similarity while diminishing negative similarities.
Masked Language Modeling (MLM)
@@
MLM is a widely used task in NLP [taylor1953mlm]. It was first introduced and applied in the BERT model [devlin2019bert]. MLM involves randomly selecting a percentage of tokens within textual data and replacing them with a special token, often denoted as MASK. The model predicts these masked tokens by taking into account the context on both sides of them, allowing the model to grasp nuanced contextual information. VLMs such as UNITER [chen2019uniter] and VisualBERT [li2019visualbert] leverage MLM for pre-training.
Masked Image Modeling (MIM)
Extending the idea of MLM to images gave rise to MIM [xie2022simmim]. In MIM, certain patches are masked, prompting the model to predict the contents of masked regions. This process enables the model to draw context from the entirety of the image, encouraging the integration of both local and global visual features. VLMs like UNITER [chen2019uniter] and ViLBERT [lu2019vilbert] leverage MIM for enhanced performance. The cross-entropy loss is employed in MLM and MIM tasks to measure the difference between predicted and actual probability distributions for the masked elements. Additionally, MLM can be combined with MIM, allowing the reconstruction of the masked signal in one modality with support from another modality [kwon2023maskedvlm].
Image-Text Matching (ITM)
ITM is another common vision-language pre-training task. Throughout the training, the model learns to map images and corresponding textual descriptions into a shared semantic space, where closely aligned vectors represent similar content in both modalities. In single-stream VLMs, the special token [CLS] represents the joint representation for both modalities. In contrast, in dual-stream VLMs, the visual and textual representations of [CLS]V and [CLS]T are concatenated. This joint representation is fed into a fully-connected layer followed by the sigmoid function, predicting a score indicating match or mismatch [chen2022vlp]. Models like CLIP [radford2021clip], ALBEF (ALign the image and text representations BEfore Fusing) [li2021albef], and METER [dou2022meter] leverage ITM during pre-training.
Combining Multiple Tasks
@@
In VLM pre-training, multiple tasks are often combined in a unified framework, allowing models to grasp nuanced contextual information across modalities. The final loss function can combine contrastive loss, cross-entropy loss for masked token prediction, and other task-specific losses. This comprehensive pre-training approach equips VLMs with versatile representations for diverse downstream tasks. For example, ALBEF [li2021albef] adopts a comprehensive pre-training objective encompassing three tasks: CL, MLM, and ITM. The overall loss is then computed as the sum of these individual components.